In recent years, 80% of those accused of Driving Under the Influence (DUI) get convicted (includes plea bargains). Suppose a firm of defense lawyers advertises that they can do much better than that. We randomly sample 20 from over 500 DUI cases handled by the law firm and find 12 convictions.

  1. Use the One Proportion to analyze these data. What do you need to type into the applet?

Probability of heads/success:

Number of tosses/sample size (n):

Number of samples:

  1. [1.5 pts] Enter your values from part f. and plot the proportion of heads/successes (instead of the number of heads/successes). In order to calculate the p-value, what would you enter into the As extreme as boxes?
Answer #1

a) We have the following advertised, H0: P=0.8 vs H1: P>0.8

From the sampled experiment,

Sample size N=20, X=12 so

Probability of success p= 12/20 = 0.6

b) To calculate p value, we need the values from above ie P=0.8

N=20 and X/N ie p= 0.6

So test statistic comes to : 0.6 - 0.8 /√(0.8*0.2)/20 = -2.236

This is the Z score and the p value is calculated from this z score.